OpenStudy (anonymous):

solve the separable differential equation for u: du/dt=e^(6u+3t) use the following initial condition: u(0)=12 so u(t)=?

OpenStudy (amistre64):

if its separable; then separate :)

OpenStudy (amistre64):

e^(a+b) = e^a e^b

OpenStudy (anonymous):

is this a differential equations course?

OpenStudy (anonymous):

e^6u+e^3t like that

OpenStudy (amistre64):

not the "+" between it, that i can see

OpenStudy (anonymous):

so du/e^(6u)=e^(3t)+dt

OpenStudy (amistre64):

there should be no "+" sign in either case lol

OpenStudy (anonymous):

its just a chapter were going through for calculus 2

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{1}{e^{6u}} du=\int\limits_{}^{}e^{3t}dt\]

OpenStudy (anonymous):

so instead du/e^(6u)=e^(3t)(dt)

OpenStudy (amistre64):

\[\frac{du}{dt}=e^{6u+3t}\] \[\frac{du}{dt}=e^{6u}\ e^{3t}\] \[\int\{\frac{du}{e^{6u}}=e^{3t}dt\}\]

OpenStudy (anonymous):

so ln(abs(e^6u))=e^3t^2+c